OpenStudy (anonymous):

what are equivalent forms of 2 1/4?

OpenStudy (texaschic101):

2 1/4 = (2 x 4 + 1) / 4 = 9/4 2 1/4 = 2.25 do you mean other forms of fractions that are equivalent ?

OpenStudy (anonymous):

You can no longer simply this fraction anymore but if u were to calculate the decimal is 2.25

OpenStudy (kendricklamar2014):

Look at @texaschic101 rolling through these math problems!!! :)

OpenStudy (texaschic101):

lol...what can I say...I like math....well...not geometry, but algebra is ok

OpenStudy (kendricklamar2014):

yah we can all see lol :)